# Cameras as Rays: Pose Estimation via Ray Diffusion

### v2

@@ 2.1 Structure-from-Motion and SLAM
Both Structure-from-motion and SLAM aim to recover camera poses and scene geometry from a large set of unordered or ordered images. Classic SfM (Snavely et al., 2006) and indirect SLAM (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021) methods generally rely on finding correspondences (Lucas & Kanade, 1981) between feature points (Bay et al., 2006; Lowe, 2004) in overlapping images, which are then efficiently optimized (Schönberger & Frahm, 2016; Schönberger et al., 2016) into coherent poses using Bundle Adjustment (Triggs et al., 1999). Subsequent work has improved the quality of features (DeTone et al., 2018), correspondences (Shen et al., 2020; Yang & Ramanan, 2019; Sarlin et al., 2020), and the bundle adjustment process itself (Tang & Tan, 2019; Lindenberger et al., 2021). On the contrary, rather than minimize geometric reconstruction errors, indirect SLAM methods (Davison et al., 2007; Schops et al., 2019) generally optimize photometric errors. While the methods described in this section can achieve (sub)pixel-perfect accuracy, their reliance on dense and large image sets makes them unsuitable for sparse-view pose estimation.
@@ 2.2 Pose Estimation from Sparsely Sampled Views
Most related to our approach are methods that can predict poses given multiple images. RelPose (Zhang et al., 2022) and RelPose++ (Lin et al., 2024) use energy-based models to compose relative rotations into sets of camera poses. SparsePose (Sinha et al., 2023) learns to iteratively refine sparse camera poses given an initial estimate, while FORGE (Jiang et al., 2024) exploits synthetic data to learn camera poses. The most comparable to us is PoseDiffusion (Wang et al., 2023), which also uses a diffusion model to denoise camera poses. However, PoseDiffusion denoises the camera parameters directly, whereas we denoise camera rays which we demonstrate to be more precise.
@@ 2.3 Ray-based Camera Parameterizations
Some prior work in calibrating generic camera representations has made use of ray-based representations of cameras, mainly for lenses such as fish-eyed for which the pinhole model is not a good approximation (Kannala & Brandt, 2006). Grossberg & Nayar (2001); Dunne et al. (2010) consider the most general camera model, where each pixel projection is modeled by its ray. Even with better calibration algorithms (Schops et al., 2020), the large number of parameters in these camera models makes them difficult to accurately calibrate. Although these works also make use of ray-based camera representations, their focus is on calibration (intrinsics) and require known calibration patterns. Neural Ray Surfaces (Vasiljevic et al., 2020) also considers learning the poses of generic cameras but does so from video rather than sparse views.
@@ 4.1 Experimental Setup
Seen Categories
COLMAP (SP+SG) (Sarlin et al., 2019) 30.7 28.4 26.5 26.8 27.0 28.1 30.6
@@
Unseen Categories
COLMAP (SP+SG) (Sarlin et al., 2019) 34.5 31.8 31.0 31.7 32.7 35.0 38.5
@@
Table 1: Camera Rotation Accuracy on CO3D (@ ${}^{\circ}$). Here we report the proportion of relative camera rotations that are within 15 degrees of the ground truth.
@@
Seen Categories
COLMAP (SP+SG) (Sarlin et al., 2019) 100 34.5 23.8 18.9 15.6 14.5 15.0
@@
Unseen Categs.
COLMAP (SP+SG) (Sarlin et al., 2019) 100 36.0 25.5 20.0 17.9 17.6 19.1
@@ Appendix A Appendix
Seen Cate.
COLMAP (SP+SG) (Sarlin et al., 2019) 57.8 56.9 56.8 57.5 58.1 58.9 60.3
@@
Unseen Cate.
COLMAP (SP+SG) (Sarlin et al., 2019) 59.1 58.8 59.1 60.2 61.2 62.6 64.6
@@
Seen Cate.
COLMAP (SP+SG) (Sarlin et al., 2019) 40.1 34.1 25.9 22.4 20.6 20.5 22.0
@@
Unseen Cate.
COLMAP (SP+SG) (Sarlin et al., 2019) 41.0 36.6 29.4 25.8 25.0 26.2 28.7
@@
Seen Categories
Ray Regression @ 5 39.8 38.5 38.7 38.7 38.7 38.6 38.4
@@
Unseen Categories
Ray Regression @ 5 29.2 29.3 29.9 29.9 30.6 29.9 30.1
@@
Seen Cate.
Ray Regression @ 0.05 100.0 74.3 57.7 49.6 45.0 41.7 39.2
@@
Unseen Cate.
Ray Regression @ 0.05 100.0 64.7 46.8 38.0 33.4 29.9 28.6
@@
Seen Categories
Apple 87.6 83.6 83.6 83.2 84.1 83.6 84.8
@@
Unseen Categories
Ball 61.7 62.8 62.8 62.9 62.3 62.2 62.1
@@
Table 12: Per-category Camera Rotation Accuracy on CO3D (@ ${}^{\circ}$) for Ray Regression.
@@
Seen Categories
Apple 100.0 92.9 87.3 81.4 78.5 74.7 74.1
@@
Unseen Categories
Ball 100.0 79.7 65.0 58.8 53.1 51.0 48.5
@@
Seen Categories
Apple 92.8 91.3 91.3 92.2 91.8 91.5 91.6
@@
Unseen Categories
Ball 73.6 74.0 74.6 74.8 76.0 74.0 75.6
@@
Table 14: Per-category Camera Rotation Accuracy on CO3D (@ ${}^{\circ}$) for Ray Diffusion.
@@
Seen Categories
Apple 100.0 96.1 92.4 89.5 87.1 85.7 84.5
@@
Unseen Categories
Ball 100.0 85.9 75.1 70.4 67.1 62.7 63.5

### v3

@@ 2.1 Structure-from-Motion and SLAM
Both Structure-from-motion and SLAM aim to recover camera poses and scene geometry from a large set of unordered or ordered images. Classic SfM (Snavely et al., 2006) and indirect SLAM (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021) methods generally rely on finding correspondences (Lucas & Kanade, 1981) between feature points (Bay et al., 2006; Lowe, 2004) in overlapping images, which are then efficiently optimized (Schönberger & Frahm, 2016; Schönberger et al., 2016) into coherent poses using Bundle Adjustment (Triggs et al., 1999). Subsequent work has improved the quality of features (DeTone et al., 2018), correspondences (Shen et al., 2020; Yang & Ramanan, 2019; Sarlin et al., 2020), and the bundle adjustment process itself (Tang & Tan, 2019; Lindenberger et al., 2021). On the contrary, rather than minimize geometric reconstruction errors, direct SLAM methods (Davison et al., 2007; Schops et al., 2019) optimize photometric errors. While the methods described in this section can achieve (sub)pixel-perfect accuracy, their reliance on dense images is unsuitable for sparse-view pose estimation.
@@ 2.2 Pose Estimation from Sparsely Sampled Views
Most related to our approach are methods that can predict poses given multiple images. RelPose (Zhang et al., 2022) and RelPose++ (Lin et al., 2024) use energy-based models to compose relative rotations into sets of camera poses. SparsePose (Sinha et al., 2023) learns to iteratively refine sparse camera poses given an initial estimate, while FORGE (Jiang et al., 2024) exploits synthetic data to learn camera poses. The most comparable to us is PoseDiffusion (Wang et al., 2023), which also uses a diffusion model to denoise camera poses. However, PoseDiffusion denoises the camera parameters directly, whereas we denoise camera rays which we demonstrate to be more precise. Concurrently to our work, PF-LRM (Wang et al., 2024a) and DUSt3R (Wang et al., 2024b) predict sparse poses by predicting pixel-aligned pointclouds (as opposed to rays in our work) and using PnP to recover cameras.
@@ 2.3 Ray-based Camera Parameterizations
Prior work in calibrating generic camera representations has used ray-based representations of cameras, mainly for fish-eyed lenses for which the pinhole model is not a good approximation (Kannala & Brandt, 2006). Grossberg & Nayar (2001); Dunne et al. (2010) consider the most general camera model, where each pixel projection is modeled by its ray. Even with better algorithms (Schops et al., 2020), the large number of parameters in these camera models makes calibration difficult. Although these works also make use of ray-based camera representations, their focus is on calibration (intrinsics) and require known calibration patterns. Neural Ray Surfaces (Vasiljevic et al., 2020) considers learning the poses of generic cameras but does so from video rather than sparse views.
@@ 4.1 Experimental Setup
Seen Categories COLMAP (SP+SG) (Sarlin et al., 2019) 30.7 28.4 26.5 26.8 27.0 28.1 30.6
@@
Unseen Categories COLMAP (SP+SG) (Sarlin et al., 2019) 34.5 31.8 31.0 31.7 32.7 35.0 38.5
@@
Table 1: Camera Rotation Accuracy on CO3D (@ 15). Here we report the proportion of relative camera rotations that are within 15 degrees of the ground truth.
@@
Seen Categories COLMAP (SP+SG) (Sarlin et al., 2019) 100 34.5 23.8 18.9 15.6 14.5 15.0
@@
Unseen Categs. COLMAP (SP+SG) (Sarlin et al., 2019) 100 36.0 25.5 20.0 17.9 17.6 19.1
@@ Appendix A Appendix
Seen Cate. COLMAP (SP+SG) (Sarlin et al., 2019) 57.8 56.9 56.8 57.5 58.1 58.9 60.3
@@
Unseen Cate. COLMAP (SP+SG) (Sarlin et al., 2019) 59.1 58.8 59.1 60.2 61.2 62.6 64.6
@@
Seen Cate. COLMAP (SP+SG) (Sarlin et al., 2019) 40.1 34.1 25.9 22.4 20.6 20.5 22.0
@@
Unseen Cate. COLMAP (SP+SG) (Sarlin et al., 2019) 41.0 36.6 29.4 25.8 25.0 26.2 28.7
@@
Seen Categories Ray Regression @ 5 39.8 38.5 38.7 38.7 38.7 38.6 38.4
@@
Unseen Categories Ray Regression @ 5 29.2 29.3 29.9 29.9 30.6 29.9 30.1
@@
Seen Cate. Ray Regression @ 0.05 100.0 74.3 57.7 49.6 45.0 41.7 39.2
@@
Unseen Cate. Ray Regression @ 0.05 100.0 64.7 46.8 38.0 33.4 29.9 28.6
@@
Seen Categories Apple 87.6 83.6 83.6 83.2 84.1 83.6 84.8
@@
Unseen Categories Ball 61.7 62.8 62.8 62.9 62.3 62.2 62.1
@@
Table 12: Per-category Camera Rotation Accuracy on CO3D (@ 15) for Ray Regression.
@@
Seen Categories Apple 100.0 92.9 87.3 81.4 78.5 74.7 74.1
@@
Unseen Categories Ball 100.0 79.7 65.0 58.8 53.1 51.0 48.5
@@
Seen Categories Apple 92.8 91.3 91.3 92.2 91.8 91.5 91.6
@@
Unseen Categories Ball 73.6 74.0 74.6 74.8 76.0 74.0 75.6
@@
Table 14: Per-category Camera Rotation Accuracy on CO3D (@ 15) for Ray Diffusion.
@@
Seen Categories Apple 100.0 96.1 92.4 89.5 87.1 85.7 84.5
@@
Unseen Categories Ball 100.0 85.9 75.1 70.4 67.1 62.7 63.5